OpenStudy (anonymous):

simplify 2y^2 x 2y^2 vx^-3 x 9x^9

OpenStudy (anonymous):

\[2y^22y^2vx^-3(9x^9)?\]

OpenStudy (anonymous):

x^-3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What is your question look like then?

OpenStudy (anonymous):

http://prntscr.com/1vji3v

OpenStudy (jack1):

simplify 2y^2 x 2y^2v^-2 vx^-3 x 9x^9 \[2y^2 \times 2y^2v^{-2} vx^{-3} \times 9x^9 \]

OpenStudy (jack1):

so first sort into like terms 2y^2 x 2y^2 x 9x^9 x x^-3 x v x v^-2 \[2y^2 \times 2y^2 \times 9x^9 \times x^{-3} \times v \times v^{-2}\] next multiply out the integers \[2 \times 2 \times 9 = 36\] next add the powers of like terms \[y^2 \times y^2 = y^{2+2} = y^4\] \[x^9 \times x^{-3} = x^{9+(-3)} = x^6\] \[v \times v^{-2} = v^{1+(-2)} = v^{-1}\] now combine \[2y^2 \times 2y^2 \times 9x^9 \times x^{-3} \times v \times v^{-2} = 36 \times y^4 \times x^6 \times v^{-1}\] \[= 36y^4 x^6 v^{-1} = \frac{36y^4 x^6} {v}\]